# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v4

@@ 4.1 Cohort’s description and pre-processing image analysis
Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in (Mitchell et al. (2023)) with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and is illustrated in Fig. 2a.
@@ 4.2 Processing of two distinct images inputs
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the Global PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracted the total overlapping image. This total overlapping image was then registered on a sulcal probabilistic atlas to unveil the model’s pattern for determining the presence or absence of the PCS.
@@ 4.3 Hyper-parameter initialization
Evaluation metrics for the explanation
@@
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized again the software developed by (Hedström et al. (2023)). As a first step we manually aligned and rescaled the two global XAI explanations (total-SHAP and total-GradCam) to the global statistical features (total-Shape). The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanations served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual input brains of the cohort were aligned in the same template and the most significant variability of each class was assigned in the total-Shape. Consequently, we anticipated the classifier to classify correctly whether a MRI has a PCS or not.
4.4 Data availability
4.5 Code availability
@@ 5.1 Classifiers performance for presence or absence of paracingulate sulcus
The performance of the simple-3D-MHL network in the left hemisphere was higher (around 73.00% in all testing metrics and 74.10% in all validation metrics) than that in the right hemisphere (around 58.00% in all testing metrics and 63.10% in the validation metrics). For the simple-3D-CNN, the performance of the network in the left hemisphere was higher (around 72.90% in all testing metrics and 74.00% in all validation metrics) than that in the right hemisphere (around 56.00% in all testing metrics and 63.00% in the validation metrics). The analytical Figures and outcomes for the simple-3D-MHL and simple-3D-CNN networks are presented in supplementary material subsection 2.2 (’Additional global explainability methods and different components PCA results’; Supplementary Extended DATA Fig. 1 a.,b.).
@@ 5.2 Global explanations and their PCA component results
We extracted and evaluated the explainability results from both networks to avoid biased observations and to investigate whether there was a clear cause-and-effect relationship between the quality of explanation and prediction performance. The hypothesis was that better prediction performance could lead to better explanation. We present the results of the simple-MHL network in the main manuscript as it had slightly better performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Extended Data Fig. 2 and 3).
@@
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey-white surface inputs mainly focus on the frontal lobe (mostly inferior lateral and inferior medial), the cingulate gyrus, the temporal lobe, and occasionally the thalamus for detecting the presence or absence of the PCS. More specifically, for the detection of the presence of PCS (paracingulate sulcus; Fig. 4) in the left and right hemisphere the simple-MHL network focuses more in the frontal lobe (medial and inferior lateral), cingulate gyrus (mostly anterior), temporal lobe and sometimes thalamus. Contrary for the absence of PCS (No paracingulate sulcus; Fig. 4) in the left hemisphere the network focuses in the frontal lobe (mostly inferior medial and inferior frontal), the temporal lobe, the cingulate gyrus, occasionally the thalamus, and specifically for the PCA-GradCam, the corpus callosum. On the right hemisphere the simple-MHL network focuses in frontal lobe (medial and lateral), the temporal lobe, and the cingulate gyrus.
@@
Fig. 5.a,b displays the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network when using the sulcal skeleton inputs. The neural network shows distributed attention but still emphasizes some key regions in both hemispheres. In the left hemisphere, for the PCS class, there is focus on the superior temporal sulcus and its branches, the posterior sylvian fissure, and parts of the central and precentral sulci, with some attention on the cingulate sulcus and the medial frontal sulcus (containing the PCS) (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on the superior and inferior temporal sulci, cingulate sulcus, medial frontal sulcus (containing the PCS), and the sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere shows similar yet less specific attention than for the PCS class, with an additional focus on the sylvian fissure and insula, and the anterior cingulate sulcus (Fig. 5a.). In the right hemisphere, additional focus is shifted to the anterior cingulate sulcus, the sub-parietal sulcus, and elements of the ventricules (Fig. 5b.).
@@
To validate the pattern indicating the presence or absence of the PCS, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Extended DATA Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey-white surface inputs, although the primary regions of focus remained consistent. Based on preliminary analysis (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam), we found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
@@
For the simple-3D-CNN on grey-white surface inputs (refer to Supplementary material Extended DATA Fig. 6 a.) in the right hemisphere, highlighted the lateral inferior frontal lobe, and inferior temporal lobe in the PCS condition and the thalamus, cingulate gyrus, lateral anterior occipital lobe, and posterior temporal lobe. In the noPCS condition, the focus was on the thalamus, cingulate gyrus, the medial frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focused on the same regions: the lateral middle frontal lobe, and inferior and superior parietal lobe, and the thalamus and in the cingulate, frontal and medial parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focused on the lateral view of the posterior temporal lobe.
@@
Faithfulness ($\uparrow$) total-SHAP 0.105 / 0.20 0.092 / 0.113 0.103 / 0.102 0.073 / 0.088
total-GradCAM 0.044 / 0.09 0.143 / 0.164 0.067 / 0.085 0.119 / 0.129
@@
a Refer to caption
b Refer to caption
c Refer to caption
@@
For the simple-3D-CNN on sulcal skeleton inputs, aimed at the accurate detection of PCS within the sulcal hemisphere inputs, the pivotal sub-regions encompassed the superior temporal sulcus, inferior precentral sulcus, sylvian fissure and sub-parietal sulcus (see Supplementary Material Extended DATA Fig. 6b). Conversely, when PCS was absent, the critical sub-regions within the right hemisphere sulcal skeleton inputs encompassed the ventricule, superior temporal sulcus, internal parietal sulcus and rostral sulcus. Transitioning to the left hemisphere, the sulcal skeleton inputs underscore the significance of the superior temporal sulcus the, sylvian fissure, and the internal parietal sulcus. When PCS was not present, the important left hemisphere sulcal skeleton inputs comprised the superior temporal sulcus, ventricule, inferior precentral sulcus, internal parietal sulcus (see Supplementary Material Extended DATA Fig. 6b).
We thereafter identified the common regions between the two networks’ outputs. The overlap results of the two networks, simple-3D-CNN and simple-3D-MHL, for the presence and absence of PCS (noPCS) reveal several common regions of interest in both the right and left hemispheres. For the presence of PCS in the right hemisphere, both networks highlight the lateral superior temporal sulcus and sylvian fissure. In the left hemisphere under the PCS condition, both models emphasize the superior temporal sulcus, the sylvian fissure, and the internal parietal sulcus. For the absence of PCS (noPCS) in the left and right hemisphere, the networks overlap in highlighting the ventricule, part of superior temporal sulcus and the internal parietal sulcus.
@@ 5.5 Ablation study of our 3D explainability framework
Table 2: Performance metrics assessing the global explanation scores of faithfulness and complexity were computed across various weight combinations assigned to the global XAI methods (total-SHAP, total-GradCam) and global feature extraction (total-Shape) within the proposed 3D-Framework for evaluating the global explanations of the simple-3D-MHL network. These combinations, denoted as 851, 815, 185, 158, 518, and 581, represent fixed-order representations of the assigned weights in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations, with weight values of 0.85 represented as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For instance, 3D-Framework-851 refers to the proposed 3D framework with weight values of 0.85 in total-Shape, 0.5 in total-SHAP, and 0.1 in total-GradCam.
Ablation study of the proposed 3D-Framework
PCS / noPCS PCS / noPCS PCS / noPCS PCS / noPCS
Faithfulness 3D-Framework-851 0.223 / 0.274 0.166 / 0.197 0.188 / 0.195 0.172 / 0.182
3D-Framework-815 0.204 / 0.233 0.207 / 0.222 0.165 / 0.174 0.192 / 0.214
3D-Framework-185 0.105 / 0.122 0.118 / 0.124 0.053 / 0.133 0.113 / 0.133
3D-Framework-158 0.074 / 0.095 0.143 / 0.173 0.122 / 0.076 0.155 / 0.162
3D-Framework-518 0.115 / 0.106 0.125 / 0.152 0.144 / 0.134 0.136 / 0.142
3D-Framework-581 0.126 / 0.138 0.087 / 0.102 0.135 / 0.142 0.103 / 0.112
Complexity 3D-Framework-851 14.584 / 14.563 14.582 / 14.582 14.579 / 14.573 14.595 / 14.592
3D-Framework-815 14.594 / 14.595 14.582 / 14.587 14.582 / 14.576 14.587 / 14.587
3D-Framework-185 14.595 / 14.584 14.585 / 14.582 14.572 / 14.595 14.584 / 14.587
3D-Framework-158 14.592 / 14.583 14.587 / 14.583 14.573 / 14.594 14.593 / 14.592
3D-Framework-518 14.583 / 14.585 14.592 / 14.584 14.573 / 14.576 14.594 / 14.594
3D-Framework-581 14.592 / 14.586 14.593 / 14.592 14.573 / 14.587 14.597 / 14.592
@@
Figure 7: A representation depicting the presence and absence of the paracingulate sulcus on the left hemisphere sulcal skeleton input a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Extended DATA Fig. 5., and NA: undefined.
@@ 6 Discussion
Our prior study (Mamalakis et al. (2024)) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. We applied the proposed guidance in a binary classification task related to a sub-region of the medial surface of the brain where secondary sulci are highly reproducible related with symptoms of psychosis, specifically hallucinations. This application was of the new-pattern discovery class. The output of the explainability indicated a wide distribution of brain regions on which the predictions depend suggesting covariant development of these regions during the perinatal period (De Vareilles et al. (2023)).
Automatic classification of psychotic and control patients based on structural MRI is a challenging task (Nunes et al. (2020)). In most cases, an acceptable detection rate is around 80.0% in clinical applications. However, in highly heterogeneous and variable cohorts, a lesser performance can be acceptable; approximately 60.0-70.0% acceptance accuracy; Iniesta et al. (2018); Hosmer and Lemeshow (2000). The TOP-OSLO cohort was particularly difficult for classification tasks of psychotic and control patients using structural MRI, with an accuracy of less than 60.0% (Nunes et al. (2020)). The variability of the paracingulate region in the TOP-OSLO cohort and the heterogeneity of the dataset create a highly challenging context for the automatic binary classification task of PCS presence. In this study, an accuracy of more than 70.0% in the left hemisphere and more than 60.0% in the right hemisphere was achieved, delivering an acceptable automated 3D deep learning network (Iniesta et al. (2018); Hosmer and Lemeshow (2000)) to apply global explainability methods for new-pattern discovery (Mamalakis et al. (2024)).
@@
We developed an XAI 3D-Framework to fill the gap in the literature by providing accurate, low-complexity, three-dimensional (3D) global explanations in the neuroimaging field, where complex representational spaces demand more than basic two-dimensional interpretations. This framework was designed to offer robust and faithful global explanations for a deep learning network in the binary classification task of PCS presence. Our global explanations surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a reliable interpretation of the deep networks for this classification task.
The overall explainability outputs cover wide regions of the brain, but we can notice some repetitive patterns through the different pipelines and modalities. In particular, there is a repeat of the cingulate region, the posterior temporal region, and both the medial and inferior frontal cortices. This may reflect some neurodevelopmental intertwining of the macroscopical development of the PCS and these regions. Regarding the cingulate and medial frontal regions, this intertwining is rather self-explanatory as the PCS is located in the medial frontal region, directly adjacent to the cingulate region, and as such the developmental events leading to the formation of a PCS are very likely to affect these regions. The two other regions mostly standing out are the inferior frontal region and the posterior temporal region. In the fetus, the sulci matching these regions (namely the inferior frontal sulcus and the posterior superior temporal sulcus) have both been reported to start appearing at 26 weeks of gestational age (w GA), while the cingulate sulcus appears earlier (around 23w GA) and the “secondary cingular sulci”, which encompass the PCS, appear later, at 31w GA (Garel et al. (2001)). This may point towards a time-window which is decisive to the development of the PCS, anterior to its actual apparition.

### v5

@@ 2.5 Cohort’s description and pre-processing image analysis
Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in (Mitchell et al. (2023)) with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and is illustrated in Fig. 2a. A more analytical description of the TOP-OSLO cohort details is available in the study Mørch-Johnsen et al. (2017).
@@ 2.6 Processing of two distinct images inputs
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the average PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracted the total overlapping image. This total overlapping image was then registered on a sulcal probabilistic atlas (the ICBM 2009a Nonlinear Asymmetric atlas, Fonov et al. (2011, 2009)) to unveil the model’s pattern for determining the presence or absence of the PCS.
@@ 2.7 Hyper-parameter initialization
2.8 Evaluation metrics for the explanation
@@
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized again the software developed by (Hedström et al. (2023)). As a first step we manually aligned and rescaled the two global XAI explanations (total-SHAP and total-GradCam) to the total-Shape. The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanation served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual input brains of the cohort were aligned in the same template and the most significant variability of each class was assigned in the total-Shape. Consequently, we anticipated the classifier to classify correctly whether a MRI has a PCS or not.
@@ 3.1 Classifiers performance for presence or absence of paracingulate sulcus
The performance of the simple-3D-MHL network in the left hemisphere was higher (around 73.00% in all testing metrics and 74.10% in all validation metrics) than that in the right hemisphere (around 58.00% in all testing metrics and 63.10% in the validation metrics). For the simple-3D-CNN, the performance of the network in the left hemisphere was higher (around 72.90% in all testing metrics and 74.00% in all validation metrics) than that in the right hemisphere (around 56.00% in all testing metrics and 63.00% in the validation metrics). The analytical Figures and outcomes for the simple-3D-MHL and simple-3D-CNN networks are presented in supplementary material subsection 2.2 (’Additional global explainability methods and different components PCA results’; Supplementary Fig. 1 a.,b.).
@@ 3.2 Global explanations and their PCA component results
We extracted and evaluated the explainability results from both networks to avoid biased observations and to investigate whether there was a clear cause-and-effect relationship between the quality of explanation and prediction performance. We present the results of the simple-MHL network in the main manuscript as it had slightly better performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Fig. 2 and 3).
@@
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey-white surface inputs mainly focus on the frontal lobe (mostly inferior lateral and inferior medial), the cingulate gyrus, the temporal lobe, and occasionally the thalamus for detecting the presence or absence of the PCS. More specifically, for the detection of the presence of PCS (paracingulate sulcus; Fig. 4) in the left and right hemisphere the simple-MHL network focuses more in the frontal lobe (medial and inferior lateral), cingulate gyrus (mostly anterior), temporal lobe and sometimes thalamus. Conversley, for the absence of PCS (No paracingulate sulcus; Fig. 4) in the left hemisphere the network focuses in the frontal lobe (mostly inferior medial and inferior frontal), the temporal lobe, the cingulate gyrus, occasionally the thalamus, and specifically for the PCA-GradCam, the corpus callosum. On the right hemisphere the simple-MHL network focuses in frontal lobe (medial and lateral), the temporal lobe, and the cingulate gyrus.
@@
Fig. 5.a,b displays the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network when using the sulcal skeleton inputs. The neural network shows distributed attention but still emphasizes some key regions in both hemispheres. In the left hemisphere, for the PCS class, there is focus on the superior temporal sulcus and its branches, the posterior sylvian fissure, and parts of the central and precentral sulci, with some attention on the cingulate sulcus and the medial frontal sulcus (containing the PCS) (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on the superior and inferior temporal sulci, cingulate sulcus, medial frontal sulcus (containing the PCS), and the sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere shows similar yet less specific attention than for the PCS class, with an additional focus on the sylvian fissure and insula, and the anterior cingulate sulcus (Fig. 5a.). In the right hemisphere, additional focus is shifted to the anterior cingulate sulcus, the sub-parietal sulcus, and elements of the ventricles (Fig. 5b.).
@@
To validate the pattern indicating the presence or absence of the PCS, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey-white surface inputs, although the primary regions of focus remained consistent. We found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
@@
For the simple-3D-CNN on grey-white surface inputs (refer to Supplementary material Fig. 6 a.) in the right hemisphere, highlighted the lateral inferior frontal lobe, and inferior temporal lobe in the PCS condition and the thalamus, cingulate gyrus, lateral anterior occipital lobe, and posterior temporal lobe. In the noPCS condition, the focus was on the thalamus, cingulate gyrus, the medial frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focused on the same regions: the lateral middle frontal lobe, and inferior and superior parietal lobe, and the thalamus and in the cingulate, frontal and medial parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focused on the lateral view of the posterior temporal lobe.
@@
a [Uncaptioned image]
b [Uncaptioned image]
c [Uncaptioned image]
@@
Faithfulness ($\uparrow$) total-SHAP 0.105 / 0.200 0.092 / 0.113 0.103 / 0.102 0.073 / 0.088
total-GradCAM 0.044 / 0.090 0.143 / 0.164 0.067 / 0.085 0.119 / 0.129
@@
For the simple-3D-CNN on sulcal skeleton inputs, aimed at the accurate detection of PCS within the sulcal hemisphere inputs, the pivotal sub-regions encompassed the superior temporal sulcus, inferior precentral sulcus, sylvian fissure and sub-parietal sulcus (see Supplementary Material Fig. 6b). Conversely, when PCS was absent, the critical sub-regions within the right hemisphere sulcal skeleton inputs encompassed the ventricule, superior temporal sulcus, internal parietal sulcus and rostral sulcus. Transitioning to the left hemisphere, the sulcal skeleton inputs underscore the significance of the superior temporal sulcus the, sylvian fissure, and the internal parietal sulcus. When PCS was not present, the important left hemisphere sulcal skeleton inputs comprised the superior temporal sulcus, ventricle, inferior precentral sulcus, internal parietal sulcus (see Supplementary Material Fig. 6b).
We thereafter identified the common regions between the two networks’ outputs. The overlap results of the two networks, simple-3D-CNN and simple-3D-MHL, for the presence and absence of PCS (noPCS) reveal several common regions of interest in both the right and left hemispheres. For the presence of PCS in the right hemisphere, both networks highlight the lateral superior temporal sulcus and sylvian fissure. In the left hemisphere under the PCS condition, both models emphasize the superior temporal sulcus, the sylvian fissure, and the internal parietal sulcus. For the absence of PCS (noPCS) in the left and right hemisphere, the networks overlap in highlighting the ventricle, part of superior temporal sulcus and the internal parietal sulcus.
Table 2: Performance metrics assessing the global explanation scores of faithfulness and complexity were computed across various weight combinations assigned to the global XAI methods (total-SHAP, total-GradCam) and global feature extraction (total-Shape) within the proposed 3D-Framework for evaluating the global explanations of the simple-3D-MHL network. These combinations, denoted as 851, 815, 185, 158, 518, and 581, represent fixed-order representations of the assigned weights in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations, with weight values of 0.85 represented as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For instance, 3D-Framework-851 refers to the proposed 3D framework with weight values of 0.85 in total-Shape, 0.5 in total-SHAP, and 0.1 in total-GradCam. The results shows the Right and Left white-grey (a) and sulcal skeleton (b) images.
(a) Ablation study of the proposed 3D-Framework
XAI metrics XAI method Left white/grey Right white/grey
PCS / noPCS PCS / noPCS
Faithfulness 3D-Framework-851 0.166 / 0.197 0.172 / 0.182
3D-Framework-815 0.207 / 0.222 0.192 / 0.214
3D-Framework-185 0.118 / 0.124 0.113 / 0.133
3D-Framework-158 0.143 / 0.173 0.155 / 0.162
3D-Framework-518 0.125 / 0.152 0.136 / 0.142
3D-Framework-581 0.087 / 0.102 0.103 / 0.112
Complexity 3D-Framework-851 14.582 / 14.582 14.595 / 14.592
3D-Framework-815 14.582 / 14.587 14.587 / 14.587
3D-Framework-185 14.585 / 14.582 14.584 / 14.587
3D-Framework-158 14.587 / 14.583 14.593 / 14.592
3D-Framework-518 14.592 / 14.584 14.594 / 14.594
3D-Framework-581 14.593 / 14.592 14.597 / 14.592
(b) Ablation study of the proposed 3D-Framework
XAI metrics XAI method Left sulcal skeleton Right sulcal skeleton
PCS / noPCS PCS / noPCS
Faithfulness 3D-Framework-851 0.223/0.274 0.188/0.195
3D-Framework-815 0.204 / 0.233 0.165 / 0.174
3D-Framework-185 0.105 / 0.122 0.053 / 0.133
3D-Framework-158 0.074 / 0.095 0.122 / 0.076
3D-Framework-518 0.115 / 0.106 0.144 / 0.134
3D-Framework-581 0.126 / 0.138 0.135 / 0.142
Complexity 3D-Framework-851 14.584 / 14.563 14.579 / 14.573
3D-Framework-815 14.594 / 14.595 14.582 / 14.576
3D-Framework-185 14.595 / 14.584 14.572 / 14.595
3D-Framework-158 14.592 / 14.583 14.573 / 14.594
3D-Framework-518 14.583 / 14.585 14.573 / 14.576
3D-Framework-581 14.592 / 14.586 14.573 / 14.587
@@ 3.5 Ablation study of our 3D explainability framework
Figure 7: A representation depicting the presence and absence of the paracingulate sulcus on the left hemisphere sulcal skeleton input a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Fig. 5., and NA: undefined.
@@ 4 Discussion
Our prior study (Mamalakis et al. (2024a)) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. We applied the proposed guidance in a binary classification task related to a sub-region of the medial surface of the brain where secondary sulci are highly reproducible related with symptoms of psychosis, specifically hallucinations. This application was of the new-pattern discovery class. The output of the explainability indicated a wide distribution of brain regions on which the predictions depend suggesting covariant development of these regions during the perinatal period (De Vareilles et al. (2023)).
Automatic classification of psychotic and control patients based on structural MRI is a challenging task (Nunes et al. (2020)). In most cases, an acceptable detection rate is around 80.0% in clinical applications. However, in highly heterogeneous and variable cohorts, a lesser performance can be acceptable; approximately 60.0-70.0% acceptance accuracy; Iniesta et al. (2018); Hosmer and Lemeshow (2000). The TOP-OSLO cohort was particularly difficult for classification tasks of psychotic and control patients using structural MRI, with an accuracy of less than 60.0% (Nunes et al. (2020)). The variability of the paracingulate region in the TOP-OSLO cohort and the heterogeneity of the dataset create a highly challenging context for the automatic binary classification task of PCS presence. In this study, an accuracy of more than 70.0% in the left hemisphere and more than 60.0% in the right hemisphere was achieved, delivering an acceptable automated 3D deep learning network (Iniesta et al. (2018); Hosmer and Lemeshow (2000)) to apply global explainability methods for new-pattern discovery (Mamalakis et al. (2024a)).
@@
We developed an innovative XAI 3D-Framework to address the need for accurate, low-complexity global explanations in neuroimaging, where traditional 2D methods fall short in capturing the intricacies of 3D representational spaces. Designed for the binary classification of PCS presence, our framework provides robust, faithful global explanations that outperform GradCam and SHAP in faithfulness. Key novelties include the integration of statistical features (Shape) with reduced dimensionality information, ensuring explanations reflect both model learning and cohort-specific variability, and the combined use of GradCam and SHAP to reduce inter-method variability and enhance reliability. This multi-method framework sets a new standard for explainable AI in neuroimaging, offering actionable insights for complex tasks like cortical morphology analysis. Our global explanations surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a reliable interpretation of the deep networks for this classification task.
The overall explainability outputs cover wide regions of the brain, but we can notice some repetitive patterns through the different pipelines and modalities. In particular, there is a repeat of the cingulate region, the posterior temporal region, and both the medial and inferior frontal cortices. This may reflect some neurodevelopmental intertwining of the macroscopical development of the PCS and these regions. Regarding the cingulate and medial frontal regions, this intertwining is self-explanatory as the PCS is located in the medial frontal region, directly adjacent to the cingulate region, and as such the developmental events leading to the formation of a PCS are very likely to affect these regions. The two other notable regions are the inferior frontal region and the posterior temporal region. In the fetus, the sulci matching these regions (namely the inferior frontal sulcus and the posterior superior temporal sulcus) have both been reported to start appearing at 26 weeks of gestational age (w GA), while the cingulate sulcus appears earlier (around 23w GA) and the “secondary cingular sulci”, which encompass the PCS, appear later, at 31w GA (Garel et al. (2001)). This may point towards a time-window which is decisive to the development of the PCS, prior to its actual apparition.